I need answer A and CModern Artifacts can produce keepsakes that will be sold for $60 each. Nondepreciation fixed costs are $2,600 per year, and variable costs are $30 per unit. The initial investment of $5,000 will be depreciated straight-line over its useful life of 5 years to a final value of zero, and the discount rate is 10%. a. What is the accounting break-even level of sales if the firm pays no taxes? (Do not round intermediate calculations. Round your answer to the nearest whole number.) Answer is complete but not entirely correct Acounting break-even level of sales ,200 units b. What is the NPV break-even level of sales if the firm pays no taxes? (Do not round intermediate calculations. Round your answer to the nearest whole number.) Answer is complete and correct NPV break-even level of sales 131 units c. what is the accounting break-even level of sales if the firms tax rate is 40%? (Do not round intermediate calculations. Round your answer to the nearest whole number.) Answer is complete but not entirely correct Acounting break-even level of sales 1,200 units
Answer #1

A accounting break even level of sales

total fixed cost = 2600+ 1000 =3600 (depreciation =1000=5000/5)

= fixed cost/ contribution= 3600/0.5= $7200 ( contribution is 50% of sale price)

no.of units to be produce to break even =$7200/$60 = 120 units

OR

fixed cost/contribution= 3600/30= 120 units ( contribution= sales price - variable cost)

B. NPV break even levels of sales if the firm pays no taxes

cash flow = (0.5* sales)-$2,600

10%, 5 year annuity factor is

1/0.10-(1/0.10*(1.10)5pwr=3.79079

projects NPV equals 0

pv(cash flows)- investment=0

3.79079*((0.50*sales)-$2,600)-$5,000 =0

(1.89539*sales) -$9,856 -$5000=0

SALES= 103,856/1.89539=$54,794

sales/ sale price =$7837.96/60= 131 units

C. tax rate is 40%of profits

calculate cash flow

cash flow= (1-T)*(revenue-cash expenses)+(T*depreciation)

=0.60(*(0.5*sales)-$2600)+(0.40*$1000)

= (0.30*sales)-1560+400

=0.30*sales-1160

the annuity factor is 3.79079,

3.79079*((0.30*sales)-1160)- $5000

1.137237*sales-4,397-5000

1.137237*sales= 9397

sales=$8,263

$8,263/60= 137 units